OpenStudy (anonymous):

integral of: x sqrt. ( 25 - (x^2))

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

sub =25-u^2

OpenStudy (anonymous):

u=25-x^2

OpenStudy (anonymous):

25 - x^2 = t -2xdx = dt int egral -0.5 root(t)dt

OpenStudy (anonymous):

its relatively easy to pick the substitutions , whatever is in the bracket or has the largest power is usually what you sub